### v1

@@ Acknowledgments
We acknowledge the support for the construction and operation of the SND@LHC detector provided by the following funding agencies: CERN; the Bulgarian Ministry of Education and Science within the National Roadmap for Research Infrastructures 2020–2027 (object CERN); ANID—Millennium Program— (Chile); the Deutsche Forschungsgemeinschaft (DFG, ID 496466340); the Italian National Institute for Nuclear Physics (INFN); JSPS, MEXT, the Global COE program of Nagoya University, the Promotion and Mutual Aid Corporation for Private Schools of Japan for Japan; the National Research Foundation of Korea with grant numbers 2021R1A2C2011003, 2020R1A2C1099546, 2021R1F1A1061717, and 2022R1A2C100505; Fundação para a Ciência e a Tecnologia, FCT (Portugal), CERN/FIS-INS/0028/2021; the Swiss National Science Foundation (SNSF); TENMAK for Turkey (Grant No. 2022TENMAK(CERN) A5.H3.F2-1). M. Climesu, H. Lacker and R. Wanke are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project 496466340. We acknowledge the funding of individuals by Fundação para a Ciência e a Tecnologia, FCT (Portugal) with grant numbers CEECIND/01334/2018, CEECINST/00032/2021 and PRT/BD/153351/2021.

### v2

@@ Acknowledgments
We acknowledge the support for the construction and operation of the SND@LHC detector provided by the following funding agencies: CERN; the Bulgarian Ministry of Education and Science within the National Roadmap for Research Infrastructures 2020–2027 (object CERN); ANID—Millennium Program— (Chile); J.C. Helo Herrera and O. Soto Sandoval acknowledge support from ANID FONDECYT grant No.1241685; the Deutsche Forschungsgemeinschaft (DFG, ID 496466340); the Italian National Institute for Nuclear Physics (INFN); JSPS, MEXT, the Global COE program of Nagoya University, the Promotion and Mutual Aid Corporation for Private Schools of Japan for Japan; the National Research Foundation of Korea with grant numbers 2021R1A2C2011003, 2020R1A2C1099546, 2021R1F1A1061717, and 2022R1A2C100505; Fundação para a Ciência e a Tecnologia, FCT (Portugal), CERN/FIS-INS/0028/2021; the Swiss National Science Foundation (SNSF); TENMAK for Turkey (Grant No. 2022TENMAK(CERN) A5.H3.F2-1). M. Climesu, H. Lacker and R. Wanke are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project 496466340. We acknowledge the funding of individuals by Fundação para a Ciência e a Tecnologia, FCT (Portugal) with grant numbers CEECIND/01334/2018, CEECINST/00032/2021 and PRT/BD/153351/2021.